OpenStudy (anonymous):

Please help!!! Determine algebraically whether the function is even, odd, or neither even nor odd. f(x) = -9 x^(3) + 8x

OpenStudy (mertsj):

If the f(x) = f(-x) the function is even If the f(x)=-f(-x) the function is odd.

OpenStudy (anonymous):

so is f(-x)=9x^(3)-8x and -f(-x)=9(-x)^(3)-8(-x) ???

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

so does -f(-x) come out to be -9x^(3)+8?

jimthompson5910 (jim_thompson5910):

f(x) = -9x^3 + 8x f(-x) = -9(-x)^3 + 8(-x) f(-x) = -9(-x^3) + 8(-x) f(-x) = 9x^3 - 8x ------------------------------------------------------- f(-x) = 9x^3 - 8x -f(-x) = -(9x^3 - 8x) -f(-x) = -9x^3 + 8x

jimthompson5910 (jim_thompson5910):

So we can see that f(x) doesn't equal f(-x) but f(x) does equal -f(-x)

jimthompson5910 (jim_thompson5910):

which makes f(x) an odd function

jimthompson5910 (jim_thompson5910):

the shortcut is to notice the exponents on the polynomial if all the exponents are even, then the polynomial function is even if all the exponents are odd, then the polynomial function is odd

OpenStudy (anonymous):

oh that makes more sense! Thank you!

jimthompson5910 (jim_thompson5910):

yw